P(x) =ax²+bx+c has positive coefficients where a, b, c are in AP. If alpha and beta are integer roots, then (alpha+beta+alpha*beta) =?

Grade:11

1 Answers

Meet
6 years ago
As we know that alpha + beta = -b/a , alpha× beta = c/a and we know that (2b/a)-(c/a)=1. So the answer is( b-a)/a